At what temperature celsius scale and fahrenheight scale gives same reading ?

Grade:11

1 Answers

Khimraj
5 years ago
formula for converting temperature is
(°F − 32) × 5/9 = °C
as °F = °C = x
x – 32 = 9x/5
-32 = 4x/5
so x = -40
Hope it clears......................